# Performance of Zero-Shot Time Series Foundation Models on Cloud Data

### v1

@@ 3 Experiments
Experimental setup   We evaluate a suite of FMs on cloud time series data. We examine several of the most recent and well know FMs for time series: VisionTS (Chen et al., 2024), TTM (revision 2) (Ekambaram et al., 2024), TimesFM (Das et al., 2024), Chronos (tiny) (Ansari et al., 2024), Moirai (small) (Woo et al., 2024) and Mamba4Cast (Bhethanabhotla et al., 2024). We look at their performance on datasets constructed out of data released on Huawei Cloud function requests, as described in Appendix A.1. We name the datasets D1, D2, D3 and D4, each of which correspond to data from a unique data centre in the Huawei Cloud. We note that, none of the FMs looked at have been pretrained on this data or cloud demand data more generally. As explained in Section 2, we perform experiments using a rolling window evaluation. This is done with a context length of $520$ and for three different forecast horizons: $30,96,336$. We selected these values as they have been widely used in previous works (Ekambaram et al., 2024; Ansari et al., 2024).
@@ A.1 Construction of the Huawei Cloud Datasets
To look at how well zero-shot FMs perform on cloud time series we construct a cloud dataset out of publicly accessible data from the Huawei Cloud. [^1] Avalible at https://github.com/sir-lab/data-release This data consists of function requests from Huawei’s serverless cloud system. The data is divided into the function requests across the data centres of five regions in China. As described in more detail in Joosen et al. (2024), the data is typical of the cloud domain having large spikes and strong periodicity (Diao et al., 2024; Darlow et al., 2024; Aksu et al., 2024). From this large store of data we curate four manageable datasets from the first two and last two regions, respectively, calling them D1, D2, D3 and D4. We disregard the third region which is much smaller than the other region’s datasets [^2] E.g., region one has $2487$ functions while Region three has $237$ functions and is generally much sparser. From the regions we selected, we downsample the number of channels by selecting ones which have a large amount of activity. We ignore channels which are near-identical copies of other selected channels, have very low activity and/or are highly random, making them unsuitable for forecasting. The indices of the selected channels are shown in the selected channels column of Table 2. After channel selection, the data is cleaned by filling missing values with zeros. It is then downsampled to 5-minute intervals by summing the values within each interval. Importantly, the selection of the channels and other dataset construction decisions was finalised before studying any forecast models and the dataset was not subsequently altered in any way when studying the performance of the FMs. An overview of the statistics of each dataset is shown in Table 2.

### v2

@@ 3 Experiments
Experimental setup   We evaluate a suite of FMs on cloud time series data. We examine several of the most recent and well know FMs for time series: VisionTS (Chen et al., 2024), TTM (revision 2) (Ekambaram et al., 2024), TimesFM (Das et al., 2024), Chronos (tiny) (Ansari et al., 2024), Moirai (small) (Woo et al., 2024) and Mamba4Cast (Bhethanabhotla et al., 2024). We look at their performance on datasets constructed out of data released on Huawei Cloud function requests, [^1] Available at https://github.com/sir-lab/time-series-fm-dataset as described in Appendix A.1. We name the datasets D1, D2, D3 and D4, each of which correspond to data from a unique data centre in the Huawei Cloud. We note that, none of the FMs looked at have been pretrained on this data or cloud demand data more generally. As explained in Section 2, we perform experiments using a rolling window evaluation. This is done with a context length of $520$ and for three different forecast horizons: $30,96,336$. We selected these values as they have been widely used in previous works (Ekambaram et al., 2024; Ansari et al., 2024).
@@ A.1 Construction of the Huawei Cloud Datasets
To look at how well zero-shot FMs perform on cloud time series we construct a cloud dataset out of publicly accessible data from the Huawei Cloud. [^2] The full Huawei Cloud data release is available at https://github.com/sir-lab/data-release This data consists of function requests from Huawei’s serverless cloud system. The data is divided into the function requests across the data centres of five regions in China. As described in more detail in Joosen et al. (2024), the data is typical of the cloud domain having large spikes and strong periodicity (Diao et al., 2024; Darlow et al., 2024; Aksu et al., 2024). From this large store of data we curate four manageable datasets from the first two and last two regions, respectively, calling them D1, D2, D3 and D4. We disregard the third region which is much smaller than the other region’s datasets [^3] E.g., region one has $2487$ functions while Region three has $237$ functions and is generally much sparser. From the regions we selected, we downsample the number of channels by selecting ones which have a large amount of activity. We ignore channels which are near-identical copies of other selected channels, have very low activity and/or are highly random, making them unsuitable for forecasting. The indices of the selected channels are shown in the selected channels column of Table 2. After channel selection, the data is cleaned by filling missing values with zeros. It is then downsampled to 5-minute intervals by summing the values within each interval. Importantly, the selection of the channels and other dataset construction decisions was finalised before studying any forecast models and the dataset was not subsequently altered in any way when studying the performance of the FMs. We have released this dataset at https://github.com/sir-lab/time-series-fm-dataset and an overview of the statistics of each dataset is shown in Table 2.